How do I calculate equilibrium national income when given T? For example: C=150+0.9yD yD=.8y I=400 T=0.2y G=700 X=130 IM=0.08y
1
Expert's answer
2013-02-28T08:49:04-0500
The level of equilibrium national income Ye = C + I + G + (X - IM) ,
where C= Consumer spending, I = Investment (Gross fixed Capital
Formation), G= Government Spending, X= Exports and IM= Imports.
Disposable income ( Yd ) is calculated as the difference between personal
income (Y) and personal tax and non tax payments (T).
Y = C + I + G + (X - IM)
Yd= Y - T
Y=150+0.9(Y-T)+400+700+130-0.08Y
Y=150+0.72Y+400+700+130-0.08Y
0,36Y=1 380
Ye=3 833 (the level of equilibrium national income)